OpenStudy (anonymous):

What is the equation in standard form of the line which passes through (–2, 6) and has a slope of –1? x + y = –4 x + y = 4 x – y = –4 x – y = 4

OpenStudy (lukecrayonz):

Solve by inputting (-2, 6) into each equation. x+y=4, -2 + 6 = 4, 4=4, true. Rewrite x+y=4 y=-x+4 y=mx+b mx is slope, b is y intercept. -x=-1 slope.